OpenStudy (anonymous):

evaluate for x=2 y=3 x^2 y^-3/x^-1 y

OpenStudy (anonymous):

Parentheses? Does the expression look like\[(x^2y^{-3})/(x^{-1}y)\]

OpenStudy (anonymous):

no its suppose to be written as a fraction.

OpenStudy (anonymous):

\[(x^2y^{-3})/(x^{-1}y)=x^3/y^4\]Now just plug in x=2 and y=3

OpenStudy (anonymous):

gracias(: